# Learning Survival Models with Right-Censored Reporting Delays

### v1

@@ 6.2 Insurance Risk Modeling
Setting. To explore the potential of expanding the eligible age range for underwriting high-risk individuals, we adopted two benchmark datasets, Dialysis (Carvalho et al. 2003) and Support2 (Knaus et al. 1995), which are available via version 0.2.9 of the SurvSet library [^2] https://github.com/chl8856/SurvSet. These datasets comprise $6805$ and $9105$ individuals, respectively. The time and event indicators in these datasets were aligned to represent accident occurrences $(z_{i}^{\prime},w_{i}^{\prime})$, where $z_{i}^{\prime}$ was standardized by dividing each value by its median. In addition to the original censoring present in the datasets, synthetic censoring times $c_{i}$ were independently generated under a constant hazard rate of $1$. The observed accident time and indicator were then redefined as $z_{i}=\min\{z_{i}^{\prime},c_{i}\}$ and $w_{i}=w_{i}^{\prime}\mathbb{I}(z_{i}^{\prime}\leq c_{i})$, respectively. Synthetic reporting times $r_{i}$ were simulated under a constant hazard rate of $\lambda_{*}=5$. Let $y_{i}=\min\{z_{i}^{\prime}+w_{i}r_{i},c_{i}\}$ and $v_{i}=w_{i}^{\prime}\mathbb{I}(z_{i}^{\prime}+w_{i}r_{i}\leq c_{i})$. In the source domain, age was employed as the covariate. We denote by Young the setting in which the youngest 25% of the population is taken as the target domain, with the remaining 75% serving as the source domain. Conversely, Senior refers to the setting where the oldest 25% is taken as the target domain and the remaining 75% as the source domain. For each of the $100$ trials, the dataset from the target domain was randomly shuffled and evenly split into two subsets. One subset, administratively censored at $\tau=0.75$, was used for parameter estimation in the second stage of the two-stage procedure. The other subset, free from administrative censoring, was employed to evaluate the estimated survival model. This test set represents individuals expected to be underwritten as new policyholders at time $\tau$ after the target domain cohort has emerged. Each individual receives a unit benefit if an accident occurs, and the premium is calculated as the integral of the estimated hazard rate for accident occurrence over the observation window. The evaluation metric is defined as the total difference between paid benefits and collected premiums, normalized by the total premiums.
@@ B.1 Standard Survival Model without Administrative Censoring
$g(y,v)\equiv\begin{cases}S(y)f_{c}(y)&(v=0)\\ f(y)S_{c}(y)&(v=1).\end{cases}$
@@
$\frac{\partial S_{c}(y)}{\partial y}=-\frac{g(y,0)}{S(y)},~~~\frac{\partial S(y)}{\partial y}=-\frac{g(y,1)}{S_{c}(y)}.$
@@
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*}),$
@@
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}.$
@@ B.2 Standard Survival Model with Administrative Censoring
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S(y)f_{c}(y).$
@@
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=f(y)S_{c}(y).$
@@ Appendix C Proofs
$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$
@@
$f_{1}(0)=\int_{0}^{\infty}f_{\circ}(0,t)dt,~~~f_{2}(0)=\int_{0}^{\infty}f_{\circ}(t,t)dt.$
@@
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*}),$
@@
$\Rightarrow\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}=\frac{\partial\log g_{v=0}(y)}{\partial y}\frac{\partial S_{c}(y)}{\partial y}-\frac{\int_{0}^{y}g_{v=1}(t,y)dt}{g_{v=0}(y)}\frac{1}{S_{c}(y)}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{2}.$
@@
$g_{v=1}(0,0)=f_{1}(0)f_{2}(0)S_{c}(0)=\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,t)}{S_{c}(t)}dt\Bigr)\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(t,t)}{S_{c}(t)}dt\Bigr).$
@@
$f_{1}(t)=\frac{g_{v=1}(t,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,s)}{S_{c}(s)}ds\Bigr)^{-1},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(s,s)}{S_{c}(s)}ds\Bigr)^{-1}.$
@@
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*}).$
@@
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S_{\circ}(y)f_{c}(y).$
@@
$\frac{\partial^{2}}{\partial y\partial z}\Pr(\bar{Z}\leq z,\bar{Y}\leq y,\bar{V}=v)=\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,V=v)=f_{\circ}(z,y)S_{c}(y).$
@@
$=\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}).$
@@
$=\mathcal{I}\Bigl[q\log\frac{\prod_{i=1}^{m}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})}{q}\Bigr]+\log\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i}).$
@@
$\mathcal{E}_{t}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})\xrightarrow{p}\mathcal{E}_{t}((\bm{\alpha}_{*},\gamma),\bm{\theta}_{2}^{*}).$
@@
$=-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)dt+\frac{n-m}{\gamma}+\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(\gamma^{-1}+\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$
@@
$\eta(\gamma)\equiv\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt-\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$
@@
$~~~~+\sum_{i=1}^{m}\int_{y_{i}-\delta_{i}}^{y_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt.$
@@
$~~~~~~~~~~~~~~+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr).$
@@
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl((\delta_{i}\nu_{i})^{2}+(y_{i}\nu_{i})\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)^{2}\Bigr)\Bigr).$
@@
$=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}-\eta(\gamma).$
@@
$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\delta_{i}^{*}y_{i}\nu_{i}^{2}}{n-m}\right).$
@@ D.1 Double Censoring Times
$f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$
@@
$=f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$
@@
$\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})\equiv\sum_{i=1}^{m}\log S_{\circ}(y_{i}\mid\bm{x}_{i})+\sum_{i=m+1}^{n}\Bigl(w_{i}\log f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})+(1-w_{i})\log\int_{z_{i}}^{y_{i}}f_{\circ}(t,y_{i}\mid\bm{x}_{i})dt\Bigr).$

### v2

@@ 6.2 Insurance Risk Modeling
Setting. To explore the potential of expanding the eligible age range for underwriting high-risk individuals, we adopted two benchmark datasets, Dialysis (Carvalho et al. 2003) and Support2 (Knaus et al. 1995), which are available via version 0.2.9 of the SurvSet library [^2] https://github.com/ErikinBC/SurvSet. These datasets comprise $6805$ and $9105$ individuals, respectively. The time and event indicators in these datasets were aligned to represent accident occurrences $(z_{i}^{\prime},w_{i}^{\prime})$, where $z_{i}^{\prime}$ was standardized by dividing each value by its median. In addition to the original censoring present in the datasets, synthetic censoring times $c_{i}$ were independently generated under a constant hazard rate of $1$. The observed accident time and indicator were then redefined as $z_{i}=\min\{z_{i}^{\prime},c_{i}\}$ and $w_{i}=w_{i}^{\prime}\mathbb{I}(z_{i}^{\prime}\leq c_{i})$, respectively. Synthetic reporting times $r_{i}$ were simulated under a constant hazard rate of $\lambda_{*}=5$. Let $y_{i}=\min\{z_{i}^{\prime}+w_{i}r_{i},c_{i}\}$ and $v_{i}=w_{i}^{\prime}\mathbb{I}(z_{i}^{\prime}+w_{i}r_{i}\leq c_{i})$. In the source domain, age was employed as the covariate. We denote by Young the setting in which the youngest 25% of the population is taken as the target domain, with the remaining 75% serving as the source domain. Conversely, Senior refers to the setting where the oldest 25% is taken as the target domain and the remaining 75% as the source domain. For each of the $100$ trials, the dataset from the target domain was randomly shuffled and evenly split into two subsets. One subset, administratively censored at $\tau=0.75$, was used for parameter estimation in the second stage of the two-stage procedure. The other subset, free from administrative censoring, was employed to evaluate the estimated survival model. This test set represents individuals expected to be underwritten as new policyholders at time $\tau$ after the target domain cohort has emerged. Each individual receives a unit benefit if an accident occurs, and the premium is calculated as the integral of the estimated hazard rate for accident occurrence over the observation window. The evaluation metric is defined as the total difference between paid benefits and collected premiums, normalized by the total premiums.
@@ B.1 Standard Survival Model without Administrative Censoring
$g(y,v)\equiv\begin{cases}S(y)f_{c}(y)&(v=0)\\ f(y)S_{c}(y)&(v=1).\end{cases}$ (20)
@@
$\frac{\partial S_{c}(y)}{\partial y}=-\frac{g(y,0)}{S(y)},~~~\frac{\partial S(y)}{\partial y}=-\frac{g(y,1)}{S_{c}(y)}.$ (21)
@@
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*}),$ (22)
@@
$\leq\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\sum_{v\in\{0,1\}}\rho(\bm{x})g(y,v\mid\bm{x};\bm{\theta}^{*},\bm{\theta}_{c}^{*})\log f(y\mid\bm{x};\bm{\theta}^{*})^{v}S(y\mid\bm{x};\bm{\theta}^{*})^{1-v}.$ (23)
@@ B.2 Standard Survival Model with Administrative Censoring
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S(y)f_{c}(y).$ (25)
@@
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=f(y)S_{c}(y).$ (26)
@@ Appendix C Proofs
$f_{1}(t)=\frac{f_{\circ}(t,t)}{f_{2}(0)},~~~f_{2}(t)=\frac{f_{\circ}(0,t)}{f_{1}(0)}.$ (30)
@@
$f_{1}(0)=\int_{0}^{\infty}f_{\circ}(0,t)dt,~~~f_{2}(0)=\int_{0}^{\infty}f_{\circ}(t,t)dt.$ (31)
@@
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log f_{\circ}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*}),$ (32)
@@
$\Rightarrow\frac{\partial^{2}S_{c}(y)}{\partial y^{2}}=\frac{\partial\log g_{v=0}(y)}{\partial y}\frac{\partial S_{c}(y)}{\partial y}-\frac{\int_{0}^{y}g_{v=1}(t,y)dt}{g_{v=0}(y)}\frac{1}{S_{c}(y)}\Bigl(\frac{\partial S_{c}(y)}{\partial y}\Bigr)^{2}.$ (34)
@@
$g_{v=1}(0,0)=f_{1}(0)f_{2}(0)S_{c}(0)=\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,t)}{S_{c}(t)}dt\Bigr)\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(t,t)}{S_{c}(t)}dt\Bigr).$ (36)
@@
$f_{1}(t)=\frac{g_{v=1}(t,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(0,s)}{S_{c}(s)}ds\Bigr)^{-1},~~~f_{2}(t)=\frac{g_{v=1}(0,t)}{S_{c}(t)}\Bigl(\int_{0}^{\infty}\frac{g_{v=1}(s,s)}{S_{c}(s)}ds\Bigr)^{-1}.$ (37)
@@
$~~~~+\int_{\mathcal{X}}d\bm{x}\int_{0}^{\infty}dy\int_{0}^{y}dz\rho(\bm{x})g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*})\log g_{v=1}(z,y\mid\bm{x};\bm{\theta}_{1}^{*},\bm{\theta}_{2}^{*},\bm{\theta}_{c}^{*}).$ (38)
@@
$\frac{\partial}{\partial y}\Pr(\bar{Y}\leq y,\bar{V}=v)=\frac{\partial}{\partial y}\Pr(Y\leq y,V=v)=S_{\circ}(y)f_{c}(y).$ (40)
@@
$\frac{\partial^{2}}{\partial y\partial z}\Pr(\bar{Z}\leq z,\bar{Y}\leq y,\bar{V}=v)=\frac{\partial^{2}}{\partial y\partial z}\Pr(Z\leq z,Y\leq y,V=v)=f_{\circ}(z,y)S_{c}(y).$ (41)
@@
$=\mathcal{E}(\bm{\theta}_{1},\bm{\theta}_{2}).$ (44)
@@
$=\mathcal{I}\Bigl[q\log\frac{\prod_{i=1}^{m}f_{1}(z_{i}\mid\bm{x}_{i})^{w_{i}}S_{1}(z_{i}\mid\bm{x}_{i})^{1-w_{i}}S_{2}(y_{i}-z_{i}\mid\bm{x}_{i})}{q}\Bigr]+\log\prod_{i=m+1}^{n}f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i}).$ (45)
@@
$\mathcal{E}_{t}((\hat{\bm{\alpha}},\gamma),\hat{\bm{\theta}}_{2})\xrightarrow{p}\mathcal{E}_{t}((\bm{\alpha}_{*},\gamma),\bm{\theta}_{2}^{*}).$ (48)
@@
$=-\sum_{i=1}^{n}\int_{0}^{y_{i}^{\dagger}}h_{b}(t)dt+\frac{n-m}{\gamma}+\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(\gamma^{-1}+\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$ (49)
@@
$\eta(\gamma)\equiv\sum_{i=1}^{m}\int_{0}^{y_{i}}h_{b}(t)S_{2}(y_{i}-t\mid\bm{x}_{i})dt-\sum_{i=1}^{m}\frac{\int_{0}^{y_{i}}\bigl(1+\gamma\int_{t}^{y_{i}}h_{b}(s)ds\bigr)h_{b}(t)\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}{1+\int_{0}^{y_{i}}h_{b}(t)\gamma\exp\bigl(\int_{t}^{y_{i}}h_{b}(s)\gamma ds\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt}.$ (50)
@@
$~~~~+\sum_{i=1}^{m}\int_{y_{i}-\delta_{i}}^{y_{i}}\Bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}+\int_{0}^{y_{i}}\nu_{i}\hat{\gamma}^{-1}S_{2}(y_{i}-s\mid\bm{x}_{i})ds\Bigr)\nu_{i}\hat{\gamma}^{-1}\gamma\exp\bigl((y_{i}-t)\nu_{i}\hat{\gamma}^{-1}\gamma\bigr)S_{2}(y_{i}-t\mid\bm{x}_{i})dt.$ (51)
@@
$~~~~~~~~~~~~~~+\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\bigr)\bigl(\epsilon_{i}(\delta_{i})+\delta_{i}\exp\bigl(\delta_{i}\nu_{i}\bigr)\bigr)\nu_{i}^{2}\Bigr).$ (52)
@@
$=\mathcal{O}\Bigl(\sum_{i=1}^{m}\Bigl((\delta_{i}\nu_{i})^{2}+(y_{i}\nu_{i})\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)+\bigl(\epsilon_{i}(\delta_{i})\nu_{i}\bigr)^{2}\Bigr)\Bigr).$ (53)
@@
$=\frac{n-m}{\gamma}-\frac{n-m}{\check{\gamma}}-\eta(\gamma).$ (54)
@@
$=1+\mathcal{O}\left(\frac{\sum_{i=1}^{m}\delta_{i}^{*}y_{i}\nu_{i}^{2}}{n-m}\right).$ (55)
@@ D.1 Double Censoring Times
$f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$ (58)
@@
$=f_{\circ}(z,y)\int_{z}^{\infty}dc_{1}\int_{\max\{y-c_{1},0\}}^{\infty}dc_{2}f_{c_{1},c_{2}}(c_{1},c_{2}).$ (61)
@@
$\tilde{\mathcal{E}}(\bm{\theta}_{1},\bm{\theta}_{2})\equiv\sum_{i=1}^{m}\log S_{\circ}(y_{i}\mid\bm{x}_{i})+\sum_{i=m+1}^{n}\Bigl(w_{i}\log f_{\circ}(z_{i},y_{i}\mid\bm{x}_{i})+(1-w_{i})\log\int_{z_{i}}^{y_{i}}f_{\circ}(t,y_{i}\mid\bm{x}_{i})dt\Bigr).$ (62)